I need your help. Can you do this for me ?Write the equation of the line that passes through (2 5) and perpendicular to the line 2y-3x=5. Write the answer in standard form
Answer by lwsshak3(11628)

Write the equation of the line that passes through (2 5) and perpendicular to the line 2y-3x=5. Write the answer in standard form.
***
equation of a line: y=mx+b, m=slope, b=y-intercept
given line:
2y-3x=5
2y=3x+5
y=(3-2)x+5/2
slope,m=3/2
..
slope of line perpendicular to given line=-2/3 (negative reciprocal)
Equation:
y=-(2/3)x+b
solve for b using coordinates of given point(2,5) on the line
5=-(2/3)*2+b
b=5+4/3=19/3
equation:
y=-(2/3)x+19/3
in std form:
2x+3y=19